OpenStudy (anonymous):

(Please help??) Find the measure of for which WXYZ is a parallelogram. A.30° B.40° C.50° D.60°

OpenStudy (anonymous):



OpenStudy (mayankdevnani):

we have to find what?

OpenStudy (anonymous):

the m<W.

OpenStudy (mayankdevnani):

okk

OpenStudy (mayankdevnani):

hint: opposite angles of parallelogram are equal

OpenStudy (anonymous):

Does it go like this? 3t-15=2t+10 or does it go.. 3t-15=90 2t+10=90 ??

OpenStudy (anonymous):

3t-15=2t+10 This one.

OpenStudy (mayankdevnani):

yeah!!! and solve it for t

OpenStudy (anonymous):

3(?)-15=2(?)+10 i dont know what number to put as T

OpenStudy (anonymous):

or is it.. 3t-15=2t+10 +15 +15 _____________________ 3t=2t+25

OpenStudy (mayankdevnani):

\[\huge 3t-15=2t+10\] \[\huge 3t-2t=10+15\] \[\huge t=25\]

OpenStudy (anonymous):

ohh..ok

OpenStudy (mayankdevnani):

and put t=25 in 3t-15 and solve

OpenStudy (anonymous):

3(25)-2(25)=10+15 ?

OpenStudy (anonymous):

25=25 ?

OpenStudy (anonymous):

3*25 - 15 75 - 15 60 degrees

OpenStudy (anonymous):

thanks. i didnt kno how to get that.

OpenStudy (mayankdevnani):

\[\huge 3t-15=\angle~m\] \[\huge 3(25)-15\] \[\huge 75-15\] \[\huge 60 ~degree=\angle~m\]